*PROJECT MANAGEMENT TEACHER NEEDED*
The President of Blue Life Company Limited, Ms. Shawna Lee is trying to decide between Projects A, B, and C to invest.

Project A requires an initial investment of $800,000 and will generate cash savings of $250,000 each year

Project B requires an initial investment of $120,000 and will generate cash savings of $30,000 each year

Project C requires an initial investment of $150,000 and will generate cash savings of $35,000 each year

N.B. Each project has an expected life of 5 years.

a) Given the initial cost and annual cash flow information what is the payback period for each project?

b) Which project should Ms Lee invest in and why?
Answer #1

a)

At payback period, sum of inflow and out flow will be 0. Let assume, t will be the payback period-

Project A -

-800000 + 250000*t = 0

250000t = 800000

t = 800000/250000 = 3.2 years

Project B -

-120000 + 30000t = 0

t = 120000/30000 = 4 years

Project C -

-150000 + 35000t = 0

t = 150000/35000 = 4.29 years

b)

From above, since Project A has the minimum payback period, hence Ms. Lee will invest in Project A.